The seeing monitor down for the night so I used the RINGSS page to report seeing. I obtained images with bad readouts in the followig images (690,719,723,724, etc). After it occurred 4x, I stopped noting which were bad. The telescope focus didn't change for one of the focus sequencessequence; Alberto had to reset M2.
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2023-04-05 | 0.0 | 10.5 | 0.0 | 0.0 | 0.0 | 10.5 |
Links
[1] mailto:sean.points@noirlab.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42350/delete?destination=print/content/enr-ctio-blanco-4-m-2023-04-05
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42350/edit?destination=print/content/enr-ctio-blanco-4-m-2023-04-05
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/20257?destination=print/content/enr-ctio-blanco-4-m-2023-04-05
Took data to determine the focus offset of other filters with respect to J. The seeing seemed to be variable when attempting this. Prelimenary dataindicate that J and H are roughly parfocal and K is J+50 units. It also seems that the NB filters 1644 [FeII], 2124 (H_2), and 2168 (Br-gamma) are parfocal with J. The focus values are close at the very least. Took calibration data of Persson 9147 in all detectors with JHKs. Took data of CD-51 9828 in 1644, 2124, and 2148 for calibration purposes.